Water of mass 10 kg was poured into the iron pot which mass is 10 kg. Calculate the quantity of heat needed to change the temperature of pot with the water in it from 10 to 100 degrees celsius.
1
Expert's answer
2020-05-27T10:32:57-0400

The amount of heat needed to change the temperature of the pot with water is "Q = c_w m_w \Delta T + c_i m_i \Delta T", where "\Delta T" is the change in temperature, "(m_w, c_w), (m_i, c_i)" are specific heat capacities and masses of water and iron respectively.

Specific heat capacities are table values: "c_w = 4200 \frac{J}{kg \cdot K}" and "c_i = 450 \frac{J}{kg \cdot K}".

Substituting "\Delta T = 90 K", "m_i = 10 kg, m_w = 10 kg" into above formula for "Q", obtain "Q \approx 4.2 \cdot 10^6 J".